OpenStudy (anonymous):

A line passes through the point (6,1) and has a slope of - 5/2 . Write an equation in slope-intercept form for this line.

OpenStudy (solomonzelman):

you know that the slope is - 5/2, so you can say y=- 5/2x+b to find b, plug in your point. y=- 5/2x+b ---> (1)=-5/2 (6) + b ---> 1=-15+b ---> b=14 so y=- 5/2x+14

OpenStudy (nincompoop):

use the point-slope form y-y1 = m(x-x1) m is your slope and x1 and y1 are the coordinates of the point given

OpenStudy (anonymous):

Thanks

OpenStudy (nincompoop):

@SolomonZelman may I ask that you do not give the answers out? it defeats the purpose of having to help people help themselves understand the concept

OpenStudy (anonymous):

both of you